OpenStudy (anonymous):

What is the axis of symmetry of the function below? f(x) = 3x2 - 24x + 19 A. x = 3 B. x = -4 C. x = -3 D. x = 4

OpenStudy (bibby):

the axis of symmetry is a vertical line with the equation of x = -b/2a.

OpenStudy (nirmalnema):

x=-24/6 x=-4